OpenStudy (anonymous):

How do I graph the following using a graphing calculator: (x+1) ^2 + (y-2)^2=9

OpenStudy (anonymous):

if you need y = something (y-2)^2 = 9 -(x+1)^2 (y-2) = +/-sqrt(9 - (x+1)^2) y = +/-sqrt(9-(x+1)^2) + 2

OpenStudy (anonymous):

I got that too. What would the graph look like? Having a hard time with my calculator.

OpenStudy (anonymous):

circle

OpenStudy (anonymous):

centered at -1,2 radius 3

OpenStudy (anonymous):

I know it's a circle but I'm not getting one on my calculator.

OpenStudy (anonymous):

\[y = 2 +/-\sqrt(9-(x+1)^2)\]

OpenStudy (anonymous):

the square root is above -(x+1)^2 as well

OpenStudy (anonymous):



OpenStudy (anonymous):

this is what i get when i type in all that info in my calc... is that what u get too?

OpenStudy (anonymous):

let x be negative as well .. the circle should look like : 

OpenStudy (anonymous):

let x be between -10 ,10 and y between -10,10 as well

OpenStudy (anonymous):

why dont i get the whole circle?

OpenStudy (anonymous):

it's two graphs .. one positive and one negative

OpenStudy (anonymous):

i put in both but it only shows me the positive side

OpenStudy (anonymous):

what is the range of values for your x and y ?

OpenStudy (anonymous):

i dont know what that is? are you talking about "window" because i dont know how that works

OpenStudy (anonymous):

*that is.

OpenStudy (anonymous):

yes the window ..

OpenStudy (anonymous):

i dont know what to put in there and how i know what to put in

OpenStudy (anonymous):

:(